OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (14, -2) and is perpendicular to 7x – 10y = 18

OpenStudy (anonymous):

so perpendicular lines : m1 * m2 = -1

OpenStudy (anonymous):

can you find the slope of the requested line now ?

OpenStudy (anonymous):

To be honest, I am really lost on this problem. I have tried to work it a bunch of different ways, but i cant figure it out, and this is a practice question, so I know my test will be hard

OpenStudy (anonymous):

ok.. so first find the slope of 7x – 10y = 18

OpenStudy (anonymous):

call it m1.. then you know that your line (lets call it's slope m2) will satisfy m1 * m2 = -1 (since they are perpendicular)